# Inexact Moreau Envelope Lagrangian Method for Non-Convex Constrained Optimization under Local Error Bound Conditions on Constraint Functions

### v1

@@ Abstract
In this paper, we study the inexact Moreau envelope Lagrangian (iMELa) method for solving smooth non-convex optimization problems over a simple polytope with additional convex inequality constraints. By incorporating a proximal term into the traditional Lagrangian function, the iMELa method approximately solves a convex optimization subproblem over the polyhedral set at each main iteration. Under the assumption of a local error bound condition for subsets of the feasible set defined by subsets of the constraints, we establish that the iMELa method can find an $\epsilon$-Karush-Kuhn-Tucker point with $\tilde{O}(\epsilon^{-2})$ gradient oracle complexity.
Keywords: non-convex constrained optimization, Lagrangian method, first-order methods, error bound conditions
@@ 1 Introduction
$f^{*}\equiv\min_{{\mathbf{x}}\in\mathcal{X}}f({\mathbf{x}})\quad\text{s.t.}\quad{\mathbf{g}}({\mathbf{x}})=(g_{1}({\mathbf{x}}),\dots,g_{m}({\mathbf{x}}))^{\top}\leq{\mathbf{0}},$ (1)
where $\mathcal{X}\subset\mathbb{R}^{n}$ is a bounded polytope that allows a computationally easy projection operator, $f$ is smooth but not necessarily convex, and $g_{i}$ for $i=1,\dots,m$ are smooth and convex. Since (1) is non-convex, computing its optimal or $\epsilon$-optimal solution is in general intractable. Therefore, the objective of this study is to compute an $\epsilon$-Karush-Kuhn-Tucker ($\epsilon$-KKT) point of (1) (see Definition 3.2) using an FOM.
@@
Under a Slater’s condition and a local error bound condition on the constraint set (Assumption 3.3), we show that the iMELa method can find an $\epsilon$-KKT point for (1) with complexity $\tilde{O}(\epsilon^{-2})$. [^1] Here and in the rest of the paper, any logarithmic factor of $\epsilon$ is omitted in notation $\tilde{O}(\cdot)$. This complexity matches the optimal complexity of an unconstrained problem [4, 5] up to a logarithmic factor and thus is also (nearly) optimal for a constrained problem. The algorithmic design and the complexity analysis of the iMELa method are motivated by [50, 51, 52], where the authors studied similar algorithms that also have complexity $O(\epsilon^{-2})$ but under different assumptions. See Section 2 for comparisons with those works in details.
Besides the complexity result, this work contributes to understanding the interaction between the convergence property of the iMELa method and the local error bound conditions of the constraints. Error bound condition has a long history of study for its connection to constraint qualifications and its impacts on the convergence of optimization algorithms (see, e.g. [24, 31, 37, 25]). Unlike existing works, where the complexity analysis relies on different constraint qualifications, e.g. Mangasarian-Fromovitz constraint qualification (MFCQ) [17], linear independence constraint qualification (LICQ) [39] and constant rank constraint qualification (CRCQ) [52], our analysis leverages the local error bound property which is weaker than most constraint qualification conditions [25]. As a result, the complexity result in this work applies to the cases which are not covered by the existing methods.
@@ 2 Related Works
In the recent literature [3, 32, 17, 16, 29], subgradient-based methods have been developed for non-smooth continuous optimization problems with convex or weakly convex inequality constraints and achieved the oracle complexity of $O(\epsilon^{-4})$ for finding a nearly $\epsilon$-KKT point. Although those methods can be applied to (1), their analysis does not utilize the smoothness of the functions and thus has a complexity higher than the one in this work when the constraints are convex.
@@
$\min_{{\mathbf{x}}}f({\mathbf{x}})+r({\mathbf{x}})\;\text{s.t.}\;{\mathbf{g}}({\mathbf{x}})=(g_{1}({\mathbf{x}}),\dots,g_{m}({\mathbf{x}}))^{\top}\leq{\mathbf{0}},\;{\mathbf{h}}({\mathbf{x}})=(h_{1}({\mathbf{x}}),\dots,h_{l}({\mathbf{x}}))^{\top}={\mathbf{0}},$ (2)
@@
where $f$, $g_{i}$, $i=1,\dots,m$, and $h_{j}$, $j=1,\dots,l$, are smooth and have Lipschitz continuous gradients, and $r$ is a proper convex lower semi-continuous function that allows a computationally easy proximal mapping. When (2) is convex, the oracle complexity of the ALM for finding an $\epsilon$-optimal solution is well studied. See, e.g. [23, 48, 47, 11, 30] for the results under different settings. We next focus on the works that are applicable to (2) or its special cases when the problem is non-convex.
@@
In several studies on ALM (e.g. [15, 10, 49]), the efficiency of an algorithm is characterized by its iteration complexity, defined as the number of main iterations needed for computing an $\epsilon$-KKT point. Since a (proximal) augmented Lagrangian subproblem must be solved in each main iteration of ALM—typically using a separate FOM—the oracle complexity is generally higher than the iteration complexity.
@@
Methods for linear constraints. Suppose $g_{i}$’s are not present and $h_{j}$’s are linear in (2). When $r\equiv 0$, Hong [15] introduced a proximal primal-dual algorithm (prox-PDA) that finds an $\epsilon$-KKT point of (2) with an iteration complexity of $O(\epsilon^{-2})$. When $r$ in (2) is the characteristic function of a box or a bounded polytope, it has been shown in [50, 51] that a smoothed proximal ALM (SP-ALM), which generalizes the classical proximal ALM [44], can achieve an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-2})$. When $r$ is the characteristic function of a compact set defined by convex inequalities, it has been shown that SP-ALM also achieves an $O(\epsilon^{-2})$ oracle complexity under CRCQ [52].
@@
Suppose no additional structural assumption is made on $r$. Hajinezhad and Hong [10] developed a variant of prox-PDA with an iteration complexity of $O(\epsilon^{-4})$. Zeng et al. [49] proposed the Moreau envelope ALM (MEAL), a type of proximal ALM, which achieves an iteration complexity of $o(\epsilon^{-2})$ when $f$ satisfies an implicit Lipschitz subgradient property and $O(\epsilon^{-2})$ when $f$ satisfies an implicit bounded subgradient property. The convergence property of MEAL is also established when the augmented Lagrangian function satisfies the Kurdyka-Łojasiewicz property [49]. Kong et al. [19, 20] proposed a quadratic penalty accelerated inexact proximal point (QP-AIPP) method, which finds an $\epsilon$-KKT point with an oracle complexity of $\tilde{O}(\epsilon^{-3})$. Under additional mildly strong assumptions, a reduced oracle complexity of $\tilde{O}(\epsilon^{-2.5})$ is achieved by an inexact ALM (iALM) by Li et al. [27], an inexact proximal accelerated augmented Lagrangian (IPAAL) by Melo et al. [33], and an inner accelerated inexact proximal augmented Lagrangian (IAIPAL) method by Kong et al. [22].
@@
Methods for nonlinear convex constraints. Suppose $g_{i}$’s are convex and $h_{j}$’s are linear in (2). Lin et al. [28] proposed an inexact proximal point penalty (iPPP) method that achieves an oracle complexity of $\tilde{O}(\epsilon^{-2.5})$ under a Slater’s condition. Under the similar assumptions, Li and Xu [26] extended this approach to a hybrid method combining ALM and the penalty method with the same complexity. Dahal et al. [8] provided a damped proximal ALM that achieves the same order of oracle complexity. The iALM by Li et al. [27] can also achieve the same complexity but requires an additional regularity assumption (see (3) below). The IAIPAL introduced in [22] was also extended by Kong et al. [21] to handle nonlinear convex constraints and obtains an $\tilde{O}(\epsilon^{-3})$ oracle complexity.
@@
When $r$ is the characteristic function of a compact set defined by convex inequalities, Pu et al. [39] developed a smoothed proximal Lagrangian method (SP-LM), which finds an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-2})$. In contrast, our method assumes $r$ is the characteristic function of a compact polytope and requires a complexity of $\tilde{O}(\epsilon^{-2})$. Besides the assumption on $r$, there are two key differences between their method and ours in the updating schemes. First, their method projects the dual variables onto a compact, artificially constructed set, whereas our method only needs to project the dual variables onto +$\mathbb{R}^{m}_{+}$. Second, their method only uses a single loop but ours employs a double-loop structure. More importantly, their method relies on a Slater-like condition and a local LICQ condition at all KKT points, whereas we assume a Slater’s condition and a local error bound condition on each active subset of the constraint set (see Assumption 3.3). We will show (see Proposition 3.4) that, if the local LICQ condition in [39] holds over all KKT points, our local error bound assumption also holds. However, the converse is not true. Thus, our work extends the complexity analysis of SP-LM to the case not covered in [39].
@@
Methods for nonconvex constraints. Suppose $g_{i}$’s are non-convex or $h_{j}$’s are nonlinear in (2). In this case, even finding a feasible solution for (2) is generally intractable. When $r$ is the characteristic function of a box and ${\mathbf{g}}\equiv{\mathbf{0}}$[6, 7] proposed an augmented Lagrangian trust-region method with the global asymptotic convergence guarantee. Their methods may produce an infeasible stationary point due to the non-convex constraints. In general, additional assumptions are needed to obtain a feasible stationary point. For example, assuming a (nearly) feasible solution can be easily obtained, the iPPP method [28] and a scaled dual descent alternating direction method of multipliers (SDD-ADMM) proposed by Sun and Sun [46] can find an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-4})$. Another common assumption made in literature is the regularity condition, which assumes that there exists $\zeta>0$ such that
@@
for any ${\mathbf{x}}\in\text{dom}(r)$. When (3) holds and $g_{i}$’s are not present, Sahin et al. [45] showed that iALM can find an $\epsilon$-KKT point with a complexity of $\tilde{O}(\epsilon^{-4})$. This complexity is reduced to $\tilde{O}(\epsilon^{-3})$ by the iALM [27] [^2] Although the problem considered in [27] involves only equality constraints, their results can be extended to the general problem (2), as noted in [27, Remark 6]. and the iPPP method [28] with the inequality constraints, and to $O(\epsilon^{-3})$ by the SDD-ADMM method [46] without the inequality constraints when there is only one block in ${\mathbf{x}}$. Moreover, the SDD-ADMM method [46] can achieve a complexity of $O(\epsilon^{-2})$ under stronger assumptions. Investigating whether an oracle complexity of $O(\epsilon^{-2})$ can be achieved by the iMELa method, SP-LM or SP-ALM under weaker assumptions than those in [27] is a topic for future study.
@@ 3 Preliminaries
[…]
For a closed convex set $\mathcal{B}$, we denote its normal cone at ${\mathbf{x}}$ by $\mathcal{N}_{\mathcal{B}}({\mathbf{x}})$. Let $\delta_{\mathcal{B}}({\mathbf{x}})$ be the zero-infinity characteristic function of $\mathcal{B}$, i.e., $\delta_{\mathcal{B}}({\mathbf{x}})=0$ if ${\mathbf{x}}\in\mathcal{B}$ and +$\infty$+ ∞ otherwise. Let $\text{proj}_{\mathcal{B}}(\cdot)$ be the projection mapping onto $\mathcal{B}$ and $\text{dist}({\mathbf{x}},\mathcal{A}):=\min_{{\mathbf{y}}\in\mathcal{A}}\|{\mathbf{x}}-{\mathbf{y}}\|$ for a set $\mathcal{A}$. Define $[\cdot]_{+}:=\max\{\cdot,0\}$. Denote $[m]=\{1,\dots,m\}$ and $[l]=\{1,\dots,l\}$.
We make the following assumptions on (1) throughout this paper.
@@
$\mathcal{L}({\mathbf{x}},\boldsymbol{\lambda}):=f({\mathbf{x}})+\sum_{i=1}^{m}\lambda_{i}g_{i}({\mathbf{x}}),$ (4)
where $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}$ is a vector of Lagrangian multipliers. A point ${\mathbf{x}}^{*}\in\mathcal{X}$ is called a Karush-Kuhn-Tucker (KKT) point of (1) if there exists $\boldsymbol{\lambda}^{*}$ such that
stationarity: $\leavevmode\nobreak\ -\nabla f({\mathbf{x}}^{*})-\textstyle\sum_{i=1}^{m}\lambda_{i}^{*}\cdot\nabla g_{i}({\mathbf{x}}^{*})\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{*}),$
feasibility: $\leavevmode\nobreak\ g_{i}({\mathbf{x}}^{*})\leq 0,\leavevmode\nobreak\ \lambda_{i}^{*}\geq 0,\;i\in[m],$
complementary slackness: $\leavevmode\nobreak\ \textstyle\sum_{i=1}^{m}\lambda_{i}^{*}g_{i}({\mathbf{x}}^{*})=0.$
Typically, a KKT point can only be approached by an algorithm as a limit point of its iterates. Within a finite number of iterations, an algorithm may only generate an $\epsilon$-KKT point for a tolerance $\epsilon>0$, which is defined as follows.
@@
[…]
@@
$J_{g}({\mathbf{x}}):=\{i\in[m]:g_{i}({\mathbf{x}})=0\}\text{ and }J_{A}({\mathbf{x}}):=\{j\in[l]:[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}=0\}$
be the index sets of the constraints active at ${\mathbf{x}}$. Let $\mathcal{X}^{*}$ denote the set of KKT points of (1). We assume the constraints of (1) satisfy a uniform local error bound condition near each KKT point.
Assumption 3.3 (Local Error Bound Condition).
There exist $\delta>0$ and $\gamma>0$ such that, for any ${\mathbf{x}}\in\mathcal{X}$, if $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta$ for some ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$, it holds that
$\textup{dist}^{2}({\mathbf{x}},\mathcal{S}(I_{g},I_{A}))\leq\gamma\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}([g_{i}({\mathbf{x}})]_{+})^{2}\\ +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\big{(}\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\big{)}^{2}\end{array}\right),$ (11)
for any $I_{g}\subset J_{g}({\mathbf{x}}^{*})$ and $I_{A}\subset J_{A}({\mathbf{x}}^{*})$.
Inequality (11) means the set $\mathcal{S}(I_{g},I_{A})$ defined in (8) satisfies a local linear error bound with $\delta$ and $\gamma$ uniform for all ${\mathbf{x}}$. We show that Assumption 3.3 holds when the constraints of (1) satisfy the local linear independence constraint qualification (LICQ) assumed in [39, Assumption 2.2], namely, there exists $\zeta>0$ such that, for any ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$,
@@
where $\sigma_{\min}(\cdot)$ denotes the smallest singular value. See the following proposition whose proof is provided in Appendix A.
@@
Throughout the paper, we assume that $p$ is a constant satisfying $p>L$. Following [50, 51, 52], given any ${\mathbf{z}}\in\mathcal{X}$, we define
$\qquad v({\mathbf{z}}):=\min_{{\mathbf{x}}\in\mathcal{X},\,g_{i}({\mathbf{x}})\leq 0,\,i\in[m]}\left\{f({\mathbf{x}})+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}\right\},$ (13)
$\begin{split}&\qquad{\mathbf{x}}({\mathbf{z}}):=\mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X},\,g_{i}({\mathbf{x}})\leq 0,\,i\in[m]}\left\{f({\mathbf{x}})+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}\right\},\\ &\text{and }\boldsymbol{\lambda}({\mathbf{z}})\text{ is a corresponding vectorof Lagrangian multipliers}.\end{split}$ (14)
Note that (13) is a convex optimization problem as $p>L$ and $v({\mathbf{z}})$ is differentiable by Danskin’s Theorem (see, e.g. [2, Proposition B.25]) with

### v2

@@ Abstract
In this paper, we investigate how structural properties of the constraint system impact the oracle complexity of smooth non-convex optimization problems with convex inequality constraints over a simple polytope. In particular, we show that, under a local error bound condition with exponent $d\in[1,2]$ on constraint functions, an inexact Moreau envelope Lagrangian method can attain an $\epsilon$-Karush–Kuhn–Tucker point with $\tilde{O}(\epsilon^{-2d})$ gradient oracle complexity. When $d=1$, this result matches the best-known complexity in literature up to logarithmic factors. Importantly, the assumed error bound condition with any $d\in[1,2]$ is strictly weaker than the local linear independence constraint qualification that is required to achieve the best-known complexity. Our results clarify the interplay between error bound conditions of constraints and algorithmic complexity, and extend complexity guarantees to a broader class of constrained non-convex problems.
@@ 1 Introduction
$\min_{{\mathbf{x}}\in\mathcal{X}}f({\mathbf{x}})\penalty 10000\ \penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ {\mathbf{g}}({\mathbf{x}})=(g_{1}({\mathbf{x}}),\dots,g_{m}({\mathbf{x}}))^{\top}\leq{\mathbf{0}},$ (1)
where $\mathcal{X}\subset\mathbb{R}^{n}$ is a bounded polytope that allows a computationally easy projection operator, $f$ is smooth but not necessarily convex, and $g_{i}$ for $i=1,\dots,m$ are smooth and convex. Since (1) is non-convex, computing its optimal or $\epsilon$-optimal solution is in general intractable. Therefore, the objective of this study is to compute an $\epsilon$-Karush–Kuhn–Tucker ($\epsilon$-KKT) point of (1) (see Definition 3.2) using an FOM.
@@
[…]
Beyond the complexity results, this work contributes to a deeper understanding of the interaction between the convergence behavior of an FOM and EBCs of constraint functions. Error bound conditions have a long history in optimization, due to their connections to constraint qualifications and their impact on the convergence of optimization algorithms (see, e.g., Lewis and Pang [24], Luo and Pang [32], Pang [38], Li [25]). While EBCs on objective functions have been extensively studied and exploited to accelerate the convergence of FOMs, the role of EBCs on constraint functions has received relatively limited attention. This work helps bridge this gap by showing how such conditions are utilized in the convergence analysis of an FOM for nonconvex constrained optimization.
@@ 2 Related Works
In the recent literature [3, 33, 17, 16, 30], subgradient-based methods have been developed for non-smooth continuous optimization problems with convex or weakly convex inequality constraints and achieved the oracle complexity of $O(\epsilon^{-4})$ for finding a nearly $\epsilon$-KKT point. Although those methods can be applied to (1), their analysis does not utilize the smoothness of the functions and thus yields a complexity higher than the one in this work when the constraints are convex.
@@
$\min_{{\mathbf{x}}}$ $\penalty 10000\ f({\mathbf{x}})+r({\mathbf{x}})$ (2)
s.t. $\penalty 10000\ {\mathbf{g}}({\mathbf{x}})=(g_{1}({\mathbf{x}}),\dots,g_{m}({\mathbf{x}}))^{\top}\leq{\mathbf{0}},$
$\penalty 10000\ {\mathbf{h}}({\mathbf{x}})=(h_{1}({\mathbf{x}}),\dots,h_{l}({\mathbf{x}}))^{\top}={\mathbf{0}},$
@@
where $f$, $g_{i}$, $i=1,\dots,m$, and $h_{j}$, $j=1,\dots,l$, are smooth and have Lipschitz continuous gradients, and $r$ is a proper convex lower semi-continuous function that allows a computationally easy proximal mapping. When (2) is convex, the oracle complexity of the ALM for finding an $\epsilon$-optimal solution is well studied. See, e.g. Lan and Monteiro [23], Xu [50, 49], He and Yuan [11], Liu et al. [31] for the results under different settings. We next focus on the works that are applicable to (2) or its special cases when the problem is non-convex.
@@
In several studies on ALM (e.g. Hong [14], Hajinezhad and Hong [10], Zeng et al. [52]), the efficiency of an algorithm is characterized by its iteration complexity, defined as the number of main iterations needed for computing an $\epsilon$-KKT point. Since a (proximal) augmented Lagrangian subproblem must be solved in each main iteration of ALM—typically using a separate FOM—the oracle complexity is generally higher than the iteration complexity.
@@
Methods for linear constraints. Suppose $g_{i}$’s are not present and $h_{j}$’s are linear in (2). When $r\equiv 0$, Hong [14] introduced a proximal primal-dual algorithm (prox-PDA) that finds an $\epsilon$-KKT point of (2) with an iteration complexity of $O(\epsilon^{-2})$. When $r$ in (2) is the characteristic function of a box or a bounded polytope, it has been shown in Zhang and Luo [53, 54] that a smoothed proximal ALM (SP-ALM), which generalizes the classical proximal ALM [45], can achieve an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-2})$. When $r$ is the characteristic function of a compact set defined by convex inequalities, it has been shown that SP-ALM also achieves an $O(\epsilon^{-2})$ oracle complexity under CRCQ [55].
@@
Suppose no additional structural assumption is made on $r$. Hajinezhad and Hong [10] developed a variant of prox-PDA with an iteration complexity of $O(\epsilon^{-4})$. Zeng et al. [52] proposed the Moreau envelope ALM (MEAL), a type of proximal ALM, which achieves an iteration complexity of $o(\epsilon^{-2})$ when $f$ satisfies an implicit Lipschitz subgradient property and $O(\epsilon^{-2})$ when $f$ satisfies an implicit bounded subgradient property. The convergence property of MEAL is also established when the augmented Lagrangian function satisfies the Kurdyka–Łojasiewicz property [52]. Kong et al. [19, 20] proposed a quadratic penalty accelerated inexact proximal point (QP-AIPP) method, which finds an $\epsilon$-KKT point with an oracle complexity of $\tilde{O}(\epsilon^{-3})$. Under additional mildly strong assumptions, a reduced oracle complexity of $\tilde{O}(\epsilon^{-2.5})$ is achieved by an inexact ALM (iALM) [27], an inexact proximal accelerated augmented Lagrangian (IPAAL) method [34], and an inner accelerated inexact proximal augmented Lagrangian (IAIPAL) method [22].
@@
Methods for nonlinear convex constraints. Suppose $g_{i}$’s are convex and $h_{j}$’s are linear in (2). Lin et al. [28] proposed an inexact proximal point penalty (iPPP) method that achieves an oracle complexity of $\tilde{O}(\epsilon^{-2.5})$ under a Slater’s condition. Under the similar assumptions, Li and Xu [26] extended this approach to a hybrid method combining ALM and the penalty method with the same complexity. Dahal et al. [8] provided a damped proximal ALM that achieves the same order of oracle complexity. The iALM by Li et al. [27] can also achieve the same complexity but requires an additional regularity assumption (see (3) below). The IAIPAL introduced in Kong et al. [22] was also extended by Kong et al. [21] to handle nonlinear convex constraints and obtains an $\tilde{O}(\epsilon^{-3})$ oracle complexity.
@@
When $r$ is the characteristic function of a compact set defined by convex inequalities, Pu et al. [40] developed a smoothed proximal Lagrangian method (SP-LM), which finds an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-2})$. In contrast, our method assumes $r$ is the characteristic function of a compact polytope and requires a complexity of $\tilde{O}(\epsilon^{-2})$. Besides the assumption on $r$, there are two key differences between their method and ours in the updating schemes. First, their method projects the dual variables onto a compact, artificially constructed set, whereas our method only needs to project the dual variables onto $\mathbb{R}^{m}_{+}$. Second, their method only uses a single loop but ours employs a double-loop structure. More importantly, their method relies on a Slater-like condition and a local LICQ condition at all KKT points, whereas we assume a Slater’s condition and a local error bound condition on each active subset of the constraint set (see Assumption 3.3). We will show (see Proposition 3.4) that, if the local LICQ condition in Pu et al. [40] holds over all KKT points, our local error bound assumption also holds. However, the converse is not true. Thus, our work extends the complexity analysis of SP-LM to the case not covered in Pu et al. [40].
@@
Methods for nonconvex constraints. Suppose $g_{i}$’s are non-convex or $h_{j}$’s are nonlinear in (2). In this case, even finding a feasible solution for (2) is generally intractable. When $r$ is the characteristic function of a box and ${\mathbf{g}}\equiv{\mathbf{0}}$, Curtis et al. [6, 7] proposed an augmented Lagrangian trust-region method with the global asymptotic convergence guarantee. Their methods may produce an infeasible stationary point due to the non-convex constraints. In general, additional assumptions are needed to obtain a feasible stationary point. For example, assuming a (nearly) feasible solution can be easily obtained, the iPPP method [28] and a scaled dual descent alternating direction method of multipliers (SDD-ADMM) proposed by Sun and Sun [48] can find an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-4})$. Another common assumption made in literature is the regularity condition, which assumes that there exists $\zeta>0$ such that
@@
for any ${\mathbf{x}}\in\text{dom}(r)$. When (3) holds on the constraints and $g_{i}$’s are not present, Sahin et al. [47] showed that iALM can find an $\epsilon$-KKT point with a complexity of $\tilde{O}(\epsilon^{-4})$. This complexity is reduced to $\tilde{O}(\epsilon^{-3})$ by the iALM [27] [^2] Although the problem considered in Li et al. [27] involves only equality constraints, their results can be extended to the general problem (2), as noted in Li et al. [27, Remark 6]. and the iPPP method [28] with the inequality constraints, and to $O(\epsilon^{-3})$ by the SDD-ADMM method [48] without the inequality constraints when there is only one block in $\mathbf{x}$}. Moreover, the SDD-ADMM method [48] can achieve a complexity of $O(\epsilon^{-2})$ under addtional stronger assumptions.
@@ 3 Preliminaries
For vectors, $\|\cdot\|$ denotes the Euclidean ($\ell_{2}$) norm and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. For matrices, $\|\cdot\|$ denotes the spectral norm. For a differentiable function $h$ on $\mathbb{R}^{n}$, we say $h$ is $L$-smooth ($L\geq 0$) on $\mathcal{X}$ if $\|\nabla h({\mathbf{x}})-\nabla h({\mathbf{x}}^{\prime})\|\leq L\|{\mathbf{x}}-{\mathbf{x}}^{\prime}\|$ for any $({\mathbf{x}},{\mathbf{x}}^{\prime})\in\mathcal{X}\times\mathcal{X}$, $\mu$-strongly convex ($\mu\geq 0$) on $\mathcal{X}$ if $h({\mathbf{x}})\geq h({\mathbf{x}}^{\prime})+\langle\nabla h({\mathbf{x}}^{\prime}),{\mathbf{x}}-{\mathbf{x}}^{\prime}\rangle+\frac{\mu}{2}\|{\mathbf{x}}-{\mathbf{x}}^{\prime}\|^{2}$ for any $({\mathbf{x}},{\mathbf{x}}^{\prime})\in\mathcal{X}\times\mathcal{X}$, and $\rho$-weakly convex ($\rho\geq 0$) on $\mathcal{X}$ if $h({\mathbf{x}})\geq h({\mathbf{x}}^{\prime})+\langle\nabla h({\mathbf{x}}^{\prime}),{\mathbf{x}}-{\mathbf{x}}^{\prime}\rangle-\frac{\rho}{2}\|{\mathbf{x}}-{\mathbf{x}}^{\prime}\|^{2}$ for any $({\mathbf{x}},{\mathbf{x}}^{\prime})\in\mathcal{X}\times\mathcal{X}$. Note that the $L$-smoothness of $h$ implies that $h$ is $L$-weakly convex but $h$ can have a smaller weak convexity constant.
For any ${\mathbf{x}}\in\mathbb{R}^{n}$ and $\delta>0$, we denote by $\mathbb{B}({\mathbf{x}},\delta)$ the Euclidean ball centered at $\mathbf{x}$} with radius $\delta$. For a closed convex set $\mathcal{B}$, we denote its normal cone at $\mathbf{x}$} by $\mathcal{N}_{\mathcal{B}}({\mathbf{x}})$. Let $\delta_{\mathcal{B}}({\mathbf{x}})$ be the zero-infinity characteristic function of $\mathcal{B}$, i.e., $\delta_{\mathcal{B}}({\mathbf{x}})=0$ if $\mathbf{x}$$\in\mathcal{B}$ and $\infty$ otherwise. Let $\text{proj}_{\mathcal{B}}(\cdot)$ be the projection mapping onto $\mathcal{B}$ and $\text{dist}({\mathbf{x}},\mathcal{A}):=\min_{{\mathbf{y}}\in\mathcal{A}}\|{\mathbf{x}}-{\mathbf{y}}\|$ for a nonempty closed set $\mathcal{A}$. Define $\cdot]_{+}:=\max\{\cdot,0\}$. Denote $m]=\{1,\dots,m\}$ and $l]=\{1,\dots,l\}$.
The following assumption is made throughout the paper.
@@
$\mathcal{L}({\mathbf{x}},\boldsymbol{\lambda}):=f({\mathbf{x}})+\textstyle\sum_{i=1}^{m}\lambda_{i}g_{i}({\mathbf{x}}),$ (4)
where $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}$ is a vector of Lagrangian multipliers. We consider a numerical method for finding an $\epsilon$-KKT point of (1) defined as follows.
@@
$\textup{dist}\big(-\nabla f({\mathbf{x}})-\textstyle\sum_{i=1}^{m}\lambda_{i}\cdot\nabla g_{i}({\mathbf{x}}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}})\big)\leq\epsilon,$ (5)
$\|[{\mathbf{g}}({\mathbf{x}})]_{+}\|\leq\epsilon,$ (6)
$\textstyle\sum_{i=1}^{m}|\lambda_{i}g_{i}({\mathbf{x}})|\leq\epsilon.$ (7)
@@
$J_{g}({\mathbf{x}}):=\{i\in[m]:g_{i}({\mathbf{x}})=0\},\quad J_{A}({\mathbf{x}}):=\{j\in[l]:[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}=0\}$
be the index sets of the constraints active at $\mathbf{x}$}. Let $\mathcal{X}^{*}$ denote the set of KKT points of (1). We assume the constraints of (1) satisfy a uniform local EBC near each KKT point.
Assumption 3.3 (Local EBC with exponent $d$).
There exist $\delta>0$, $\gamma>0$, and $d\in[1,2]$ such that, for any ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ and any ${\mathbf{x}}\in\mathbb{B}({\mathbf{x}}^{*},\delta)$, it holds for every index pair $I=(I_{g},I_{A})$ with $I_{g}\subseteq J_{g}({\mathbf{x}}^{*})$ and $I_{A}\subseteq J_{A}({\mathbf{x}}^{*})$ that
[…]
Inequality (14) states that each active feasible set $\mathcal{S}(I_{g},I_{A})$ defined in (10) satisfies a local EBC with exponent $d\in[1,2]$, where the constants $\delta$ and $\gamma$ are uniform over all $\mathbf{x}$} sufficiently close to $\mathcal{X}^{*}$ and all admissible index pairs $(I_{g},I_{A})$. Similar EBCs have been studied in constrained optimization; see, e.g., Yang and Lin [51], Lin et al. [29]. We further show that Assumption 3.3 holds whenever the constraints in (1) satisfy the local linear independence constraint qualification (LICQ) assumed in Pu et al. [40, Assumption 2.2]. Specifically, the local LICQ requires that there exists $\zeta>0$ such that, for any ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$,
@@
where $\sigma_{\min}(\cdot)$ denotes the smallest singular value. Under (15), each set $\mathcal{S}(I_{g},I_{A})$ satisfies a local EBC with $d=1$, which in turn implies Assumption 3.3 for any $d\in[1,2]$. On the contrary, Assumption 3.3 does not imply the local LICQ and thus is strictly weaker. See the following proposition whose formal proof is provided in Appendix A.
@@
Throughout the paper, we assume that $p$ is a constant satisfying $p>L$. Following Zhang and Luo [53, 54], Zhang et al. [55], given any ${\mathbf{z}}\in\mathcal{X}$, we define
$v({\mathbf{z}}):=$ $\penalty 10000\ \min_{{\mathbf{x}}\in\mathcal{X},\,g_{i}({\mathbf{x}})\leq 0,\,i\in[m]}\left\{f({\mathbf{x}})+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}\right\},$ (16)
${\mathbf{x}}({\mathbf{z}}):=$ $\penalty 10000\ \mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X},\,g_{i}({\mathbf{x}})\leq 0,\,i\in[m]}\left\{f({\mathbf{x}})+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}\right\},$ (17)
and $\boldsymbol{\lambda}({\mathbf{z}})$ is the optimal Lagrangian multipliers corresponding to ${\mathbf{x}}({\mathbf{z}})$. Note that (16) is a convex optimization problem as $p>L$ and $v({\mathbf{z}})$ is differentiable by Danskin’s Theorem (see, e.g. Bertsekas [2, Proposition B.25]) with